In a right triangle, one of the acute angles is fourteen times as large as the other acute angle. what are measures of the two acute angles?
Answer by checkley77(9343)

x+14x=90
15x=90
x=90/15
x=6 ans. for the smaller angle.
14*6=84 ans. for the larger angle.
Proof:
6+84=90
90=90